OpenStudy (anonymous):

factor the trinomial completely by using any method. 3wsquared-w-2

OpenStudy (vishweshshrimali5):

\[\large{3w^2 - w - 2}\] \[\large{= 3w^2 - 6w + w - 2}\] \[\large{= 3w(w-2) + 1(w-2)}\] \[\large{= (3w+1)(w-2)}\]